OpenStudy (anonymous):

I've been stuck on this problem for so long :( Find the double integral by interpreting it as the volume of a solid.

OpenStudy (anonymous):

You might want to try some of that Gorilla glue remover. They should sell it at your local walmart. Hope this helped! :D

OpenStudy (anonymous):

\[\int\limits_{}^{}\int\limits_{}^{}\sqrt{4-x ^{2}-y ^{2}}\]

OpenStudy (anonymous):

\[x ^{2}+y ^{2}\le4\]

OpenStudy (anonymous):

and x and y are both greater than or equal to 0 of course

OpenStudy (anonymous):

what does sqrt(4 - x^2 - y^2) look like

OpenStudy (anonymous):

a paraboloid

OpenStudy (anonymous):

no

OpenStudy (anonymous):

a circle?

OpenStudy (anonymous):

with radius 2?

OpenStudy (anonymous):

it's a upper hemisphere

OpenStudy (anonymous):

with radius two

OpenStudy (anonymous):

okay, so you'd compute the area of the hemisphere using 2 as the radius?

OpenStudy (anonymous):

you're supposed to interpret the double integral above using the volume of the hemisphere

OpenStudy (anonymous):

which is 2/3pir^2

OpenStudy (anonymous):

okay?

OpenStudy (anonymous):

Still confused

OpenStudy (anonymous):

2/3pi(2)^2?

OpenStudy (anonymous):

ans: "the double integral can be interpreted as the volume of the hemisphere whose radius is 2"